# Detecting Discrepancies Between AI-Generated and Natural Images Using Uncertainty

### v1

@@ 3.2 Uncertainty Definition
3.3 Uncertainty Calculation
The predictive uncertainty is typically calculated as the variance of predictions obtained with certain perturbations. In this work, we simply leverage $\theta$ as the features or parameters before perturbation. Specifically, the predictive uncertainty $u(\mathbf{x})$ can be calculated by,
@@ 3.3 Uncertainty Calculation
where $n$ is the number of samples, $f(\mathbf{x};\theta_{k})$ denotes the L2-normalized features of an input image $\mathbf{x}$ when inferring with the parameter $\theta_{k}$, and $\theta_{t}$ stands for the teacher model used in DINOv2.
@@
$u(\mathbf{x})\leq\frac{1}{n}\sum_{k=1}^{n}\left\|f(\mathbf{x};\theta_{k})-\frac{1}{n}\sum_{j}^{n}f(\mathbf{x};\theta_{j})\right\|^{2}\left\|f(\mathbf{x};\theta_{t})\right\|^{2}=2-\frac{2}{n}\sum_{k=1}^{n}f(\mathbf{x};\theta_{k})^{\top}f(\mathbf{x};\theta),$ (3)
@@
where $\theta$ denotes the parameter before injecting perturbation, and we leverage an unbiased assumption that the expectation $\mathbb{E}_{\theta_{j}}f(\mathbf{x};\theta_{j})$ approaches the feature $f(\mathbf{x};\theta)$ extracted by the non-perturbed parameter $\theta$. Eq. 3 provides a simple approach to calculate the uncertainty without needing a teacher model used in the training phase of DINOv2. The insight of Eq. 3 is intuitive. Specifically, if an image $\mathbf{x}$ causes a high feature similarity between the original and perturbed parameter, the image leads to a low uncertainty and is more likely to be a natural image.
3.4 An overview of WePe
As discussed above, the proposed WePe is based on a large model pre-trained on a large number of natural images. In this work, we chose DINOv2 (Oquab et al., 2024), a large model trained with contrastive learning on image data. In order to capture the uncertainty of the model on the test images, we extract image features using the original model and the model after adding noise to the parameters respectively. The similarity between the pre-perturbation and post-perturbation feature vectors is quantified using a suitable distance metric, such as cosine similarity. Images exhibiting high similarity are classified as real, while those with low similarity are identified as generated. This method not only capitalizes on the characteristics of the DINOv2 model but also provides a robust framework for distinguishing between real and generated images based on their feature stability under model perturbation.
3.5 Discussions
Why choose DINOv2? In addition to DINOv2, CLIP is a commonly used model. However, our experiments show that CLIP performs sub-optimally compared to DINOv2 (see Table 5). We believe this difference stems from their training strategies. Unlike DINOv2, CLIP is a multimodal model that combines image and textual features from captions for contrastive learning, which may lead it to focus on broader semantic features rather than fine details. In contrast, DINOv2 emphasizes contrastive learning solely on images, allowing it to better capture subtle differences between natural and generated images. Therefore, we use DINOv2 in our experiments.
Why choose weight perturbation? Common methods for measuring uncertainty include MC Dropout and Deep Ensembles. MC Dropout involves keeping dropout active during testing and performing multiple forward passes on the inputs to generate outputs with different network structures. The variability among these outputs serves as an estimate of the model’s uncertainty regarding the input data. However, since DINOv2 does not utilize dropout during training, MC Dropout may not yield optimal results (see Table 5). Deep Ensembles, on the other hand, trains multiple networks independently and uses the differences in their outputs on test samples to assess uncertainty. However, training multiple DINOv2-level models is impractical. Therefore, in our study, we choose to perturb the model parameters and assess the differences in outputs from the original and perturbed models to estimate uncertainty for the test images.
Why does weight perturbation work? Incorporating weight perturbations during testing can effectively simulate Bayesian inference, thereby capturing the inherent uncertainty in neural networks. From a Bayesian perspective, the weights of a neural network are not fixed but rather distributions reflecting the range of plausible values given the data. By introducing noise into the network weights at test time, we mimic the process of drawing samples from a posterior distribution over weights, a core concept in Bayesian inference. This technique enables the model to generate diverse predictions, reflecting its uncertainty, particularly when encountering out-of-distribution samples. This approach is akin to Bayesian neural networks, where weight uncertainty is explicitly modeled. For example, (Blundell et al., 2015) proposed using variational methods to approximate weight distributions in Bayesian neural networks, allowing for uncertainty quantification through weight perturbations. Similarly, (Gal & Ghahramani, 2016) demonstrated that introducing dropout at test time serves as a Bayesian approximation, with the added noise acting as a proxy for weight sampling, thus allowing for reliable uncertainty estimation. In this paper, we capture uncertainty by adding noise directly to the model parameters to distinguish between natural and generated images. However, trying to theoretically analyze the different sensitivities of natural and generated images to model weights in the representation space of a large model is a difficult task. Therefore, we only draw this conclusion through various empirical observations in this paper, and do not directly prove this theoretically. The theoretical proof will be left to our future work.
4 Experiments
4.1 Experiment setup
Datasets. Following previous work (He et al., 2024), we evaluate the performance of WePe on ImageNet (Deng et al., 2009), LSUN-BEDROOM (Yu et al., 2015) and GenImage (Zhu et al., 2023). For ImageNet and LSUN-BEDROOM, the generated images are provided by  (Stein et al., 2023). For ImageNet, the generative models include ADM (Dhariwal & Nichol, 2021), ADM-G, LDM (Rombach et al., 2022), DiT-XL2 (Peebles & Xie, 2023), BigGAN (Brock et al., 2019), GigaGAN (Kang et al., 2023), StyleGAN (Karras et al., 2019), RQ-Transformer (Lee et al., 2022), and MaskGIT (Chang et al., 2022). For LSUN-BEDROOM,the generative models include ADM, DDPM (Ho et al., 2020), iDDPM (Nichol & Dhariwal, 2021), Diffusion Projected GAN (Wang et al., 2023b), Projected GAN (Wang et al., 2023b), StyleGAN (Karras et al., 2019) and Unleasing Transformer (Bond-Taylor et al., 2022). GenImage primarily employs the Diffusion model for image generation. The fake images are generated by Stable Diffusion V1.4 (Rombach et al., 2022), Stable Diffusion V1.5 (Rombach et al., 2022), GLIDE, VQDM (Gu et al., 2022), Wukong (Wukong, ), BigGAN, ADM, and Midjourney (Midjourney, ). It contains 1,331,167 real and 1,350,000 generated images.
Evaluation metrics. Following RIGID, we mainly use the following metrics: (1) the average precision (AP), (2) the area under the receiver operating characteristic curve (AUROC). On GenImage, since the dataset is too large and re-implementing all baselines is time-consuming, we obtain the results directly from the corresponding papers and report the classification accuracy (ACC).
Baselines. Following RIGID, we take both training-free methods and training methods as baselines. For training-free methods, we take RIGID (He et al., 2024) and AEROBLADE (Ricker et al., 2024) as baselines. For training methods, we take DIRE (Wang et al., 2023a), CNNspot (Wang et al., 2020), Ojha (Ojha et al., 2023) and NPR (Tan et al., 2024) as baselines. Besides, on GenImage, we also report the result of Frank (Frank et al., 2020), Durall (Durall et al., 2020), Patchfor (Chai et al., 2020), F3Net (Qian et al., 2020), SelfBland (Shiohara & Yamasaki, 2022), GANDetection (Mandelli et al., 2022), LGrad (Tan et al., 2023), ResNet-50 (He et al., 2016), DeiT-S (Touvron et al., 2021), Swin-T (Liu et al., 2021), Spec (Zhang et al., 2019), GramNet (Liu et al., 2020). For these baselines, we get the results directly in the corresponding papers without reimplementing them.
Experiment details. To balance detection performance and efficiency, we use DINOv2 ViT-L/14. We report the average results under five different random seeds and report the variance in Figure 8. In our experiments we find that perturbing the high layers may lead to a large corruption in the features of the real images, resulting in sub-optimal results. Therefore, We do not perturb the high-level parameters. In DINOv2 ViT-L/14, the model has 24 transformer blocks, and we only perturb the parameters of the first 19 blocks with Gaussian perturbations of zero mean. The variance of the added Gaussian noise is proportional to the mean value of the parameters in each block, with the ratio set to 0.1. Ablation experiments show that our method is robust to the blocks chosen for perturbation and the level of noise.
4.2 Results
Comparison with other baselines. We conduct full comparative experiments on the three benchmarks mentioned. As shown in Table 1,  2 and  3, WePe achieves the best detection performance on ImageNet, LSUN-BEDROOM and GenImage without the need for training. It is worth noting that on the large-scale GenImage benchmark, the training-based method, despite having perfect performance on the generators used during training, performs poorly on many generators not seen during training. This illustrates the extreme dependence of the performance of training-based methods on the diversity of the training set. In contrast, our method does not require training, performs well on a wide variety of generators, and outperforms the SOTA training method by $3.36\%$ on average. On generators that have not been seen during training, such as VQDM, many training methods exhibit random prediction results, whereas WePe achieves superior detection performance. To further illustrate the effectiveness of our method, we count the difference in feature similarity between real and fake images on the pre- and post-perturbation models. As shown in Figure 3, the small perturbation of the model has less effect on the real images than on generated images, resulting in higher feature similarity before and after the perturbation. The discrepancy effectively distinguishes the real image from the generated image.
@@ 4.2 Results
Figure 3: Comparison of cosine similarity between original and transformed images. The generated images are from: (a) ADM, (b) BigGAN, and (c) DDPM.
@@
Table 1: AI-generated image detection performance on ImageNet. Values are percentages. Bold numbers are superior results and the underlined italicized values are the second-best performance. A higher value indicates better performance.

### v2

@@ 3.2 Uncertainty estimation via weight perturbation
In this study, we leverage DINOv2 for uncertainty estimation, capitalizing on two key advantages. First, DINOv2 is pre-trained on an extensive datasets, endowing it with robust generalization capabilities. This broad pre-training enables the model to effectively capture the distributional characteristics of natural images, thereby enhancing the generalizability of our proposed method across diverse scenarios. Second, as DINOv2 is exclusively pre-trained on natural images, it exhibits distinct uncertainty profiles when processing natural versus AI-generated images. This differential uncertainty provides a reliable basis for distinguishing between the two image types. DINOv2 is a self-supervised learning model, employing a student-teacher framework where the student model $\theta$ is trained to align with the representations of the teacher model $\theta_{t}$. Consequently, we utilize the feature similarity between the embeddings produced by the teacher and student models as the prediction:
$\hat{y}(x)=f(\mathbf{x};\theta)^{\top}f(\mathbf{x};\theta_{t}),$ (11)
where $f(\mathbf{x};\theta)$ denotes the L2-normalized features of an input image $\mathbf{x}$ when inferring with the parameter $\theta$. However, estimating uncertainty with DINOv2 presents challenges in obtaining multiple predictions. First, as DINOv2 does not employ dropout during training, MC Dropout may yield suboptimal results (see Table 2). Second, training multiple DINOv2-scale models is computationally infeasible, rendering Deep Ensemble impractical. Instead, we use an alternative approach, weight perturbation, to obtain multiple predictions. Specifically, let $\theta_{k}$ denote the perturbed parameters of the student network, according to Eq. 10, the predictive uncertainty $u(\mathbf{x})$ can be calculated by,
@@
$\begin{split}u(\mathbf{x})\leq\frac{1}{n}\sum_{k=1}^{n}\left\|f(\mathbf{x};\theta_{k})-\frac{1}{n}\sum_{j}^{n}f(\mathbf{x};\theta_{j})\right\|^{2}\left\|f(\mathbf{x};\theta_{t})\right\|^{2}=2-\frac{2}{n}\sum_{k=1}^{n}f(\mathbf{x};\theta_{k})^{\top}f(\mathbf{x};\theta),\end{split}$ (13)
@@
where $\theta$ denotes the parameter of student model before injecting perturbation, and we leverage an unbiased assumption that the expectation $\mathbb{E}_{\theta_{j}}f(\mathbf{x};\theta_{j})$ approaches the feature $f(\mathbf{x};\theta)$ extracted by the non-perturbed parameter $\theta$. Eq. 13 provides a simple approach to calculate the uncertainty without needing a teacher model used in the training phase of DINOv2. We provide an analysis of the validity of this upper bound in Appendix D. The insight of Eq. 13 is intuitive. Specifically, if an image $\mathbf{x}$ causes a high feature similarity between the original and perturbed parameter, the image leads to a low uncertainty and is more likely to be a natural image.
@@
Figure 4: Comparison of cosine similarity between features on original and perturbed models. The generated images are from: (a) ADM, (b) BigGAN, and (c) DDPM.
3.3 Theoretical analysis of the effectiveness of weight perturbation
In this section, we present a theoretical analysis elucidating why weight perturbations effectively distinguish between natural and AI-generated images. We establish a formal metric to quantify the sensitivity of neural network feature representations to parameter perturbations and demonstrate its differential behavior across natural images and AI-generated images.
To formalize this notion, we introduce the following definition of perturbation sensitivity:
Definition 3.1 (Perturbation Sensitivity).
For a neural network $f:\mathcal{X}\to\mathbb{R}^{d}$, parameterized by $\theta\in\mathbb{R}^{p}$, which maps an input $x\in\mathcal{X}$ to a feature vector $f(x;\theta)$, the sensitivity to parameter perturbations is defined as:
$\text{sen}(x)=\|\nabla_{\theta}f(x;\theta)\|_{F}^{2},$ (14)
where $\nabla_{\theta}f(x;\theta)$ denotes the Jacobian matrix of the feature mapping with respect to $\theta$, and $\|\cdot\|_{F}$ represents the Frobenius norm.
This metric captures the magnitude of variation in the feature representation induced by infinitesimal changes in the model parameters, providing a robust measure of sensitivity to weight perturbations.
We proceed to establish the differential sensitivity of the model across distributions through the following theorem:
Theorem 3.2 (Differential Sensitivity).
Let a neural network $f(x;\theta)$ be trained on a large amount of natural images sampled from natural image distribution $\mathcal{D}^{1}$: $T=\{(x^{1},y^{1}),(x^{2},y^{2}),...,(x^{n},y^{n})\}\sim\mathcal{D}^{1}$. The expected sensitivity of the feature representations to parameter perturbations is lower for inputs drawn from $\mathcal{D}^{1}$ compared to those drawn from a generated image distribution $\mathcal{D}^{0}$, generated by generative models. Formally:
$\mathbb{E}_{x\sim\mathcal{D}^{1}}\left[\text{sen}(x)\right]\leq\mathbb{E}_{x\sim\mathcal{D}^{0}}\left[\text{sen}(x)\right].$ (15)
This theorem asserts that feature representations of natural images, optimized through training on $\mathcal{D}^{1}$, exhibit greater robustness to parameter perturbations than those of AI-generated images, which lie outside the training distribution. This differential sensitivity underpins the efficacy of weight perturbations in AI-generated image detection, enabling the model to distinguish natural images from AI-generated images. A rigorous proof of this theorem is provided in Appendix C.
@@ 3.3 Theoretical analysis of the effectiveness of weight perturbation
Table 1: AI-generated image detection performance on ImageNet. Values are percentages. Bold numbers are superior results. We compare training methods and training-free methods separately.